Skater A passes skater B with a constant velocity of 12 m/s. Three seconds later, skater B starts after skater A with an acceleration of 4 m/s2. How long does it take skater B to catch skater A?
1
Expert's answer
2021-08-18T07:02:15-0400

Equations of motion of a skaters are given by

"x_A=36+12t\\
x_B=2t^2"

When the skater B catches skater A

"x_A=x_B\\
36+12t=2t^2"

We get

"t^2-6t-18=0""t=8.2\:\rm s"